# Deep Hedging with Options

### v2

@@ 3.1   Daily implied volatility surfaces
The time-$t$ IV of an option with time-to-maturity $\tau_{t}=\frac{T-t}{252}$ years and moneyness $M_{t}=\frac{1}{\sqrt{\tau_{t}}}\log\frac{S_{t}e^{(r_{t}-q_{t})\tau_{t}}}{K}$ is given by:
@@
The vector $\beta_{t}=(\beta_{t,1},\beta_{t,2},\beta_{t,3},\beta_{t,4},\beta_{t,5})$ represents the IV factor coefficients at time $t$, while the functions $\{f_{i}\}_{i=1}^{5}$ allow representing the long-term at-the-money (ATM) level, the time-to-maturity slope, the moneyness slope, the smile attenuation, and the smirk, respectively. A detailed description of the functional components $\{f_{i}\}_{i=1}^{5}$ of the IV surface can be found in Appendix B.1.
@@ 3.2   Joint implied volatility and return
The JIVR model introduced by François et al., (2023) builds upon the IV representation in Equation (10), offering an explicit formulation for the joint dynamics of the IV surface and the S&P 500 price. More precisely, this joint representation is based on an econometric model for (i) the underlying asset returns, and (ii) fluctuations of the IV surface coefficients $\beta_{t}$ along with a mean-reversion component for their volatilities $h_{t}$. The multivariate time series formulation of the JIVR model is provided in detail in Appendix B.2.
The JIVR model is used in subsequent simulation experiments to generate paths of the state variables $(S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$, which drive the market dynamics, where $h_{t,R}$ and $\{h_{t,i}\}^{5}_{i=1}$ are volatilities for the S&P 500 and each of the IV factors. Estimates of model parameters and volatility series $\{\hat{h}_{t,i}\}_{t=1}^{N}$ with $i\in\{1,\ldots,5,R\}$ are taken from François et al., (2023), who apply maximum likelihood on a multivariate time series made of S&P 500 returns and surface coefficients estimates $\{\hat{\beta}_{t}\}_{t=1}^{N}$, with sample dates extending between January 4, 1996 and December 31, 2020.
@@ 4.1   Market settings for numerical experiments
We consider daily trading periods. Initial conditions of the JIVR model, ($\left(\{\beta_{0,i}\}^{5}_{i=1},h_{0,R},\{h_{0,i}\}^{5}_{i=1}\right)$, are randomly sampled from the daily estimated values in our data set, covering the period from January 4, 1996, to December 31, 2020. Across all experiments, the annualized continuously compounded risk-free rate and dividend yield are assumed to remain constant, with values fixed at $r=2.66\%$ and $q=1.77\%$, respectively. [^7] The annualized rates of the S&P 500 dividend yield (1.77%) and the zero-coupon yield (2.66%) are calculated as the average over the sample period from January 4, 1996, to December 31, 2020, using OptionMetrics data.
The initial value of the underlying asset is set to $S_{0}=100$. The hedged portfolio is an ATM straddle with a maturity of $T=63$ days. At any time $t<T$, the portfolio value $\mathcal{P}_{t}$ is determined using the IV surface prevailing at that moment. At maturity $\mathcal{P}_{T}$ represents the final portfolio value, which is the straddle payoff in our example.
The hedging instruments are the risk-free asset, the underlying asset, and an option with a maturity longer than that of the straddle—specifically, an ATM European call option with a maturity of $T^{*}=84$ days. The time-to-maturity of the hedging option naturally decreases over time and is not reset to 84 days at each rebalancing step. Positions in all hedging instruments are rebalanced daily.
The hedge follows the self-financing dynamics in Equation (1), incorporating proportional transaction costs on the hedging option, too. As reported in Chaudhury, (2019), the average cost for S&P 500 index call options is 0.95%. To evaluate its impact, we consider $\kappa_{2}\in\{0.5\%,1\%,1.5\%,2\%\}$. In contrast, transaction costs for the underlying asset are negligible, around 0.047% according to Bazzana and Collini, (2020). We set $\kappa_{1}=0.05\%$. The initial hedging portfolio value matches the straddle price, i.e., $V_{0}=P_{0}$.
@@ 4.2   Benchmarks
For the second and third benchmarks, the delta and gamma of financial instruments are computed using the practitioner’s approach. This involves inserting the IV for each instrument into the closed-form expressions for Black-Scholes’ delta and gamma. In the case of delta hedging, the delta is adjusted based on the correction introduced by Leland, (1985), which accounts for the impact of proportional transaction costs on the underlying asset position. This adjusted delta reverts to the standard Black-Scholes delta when no transaction costs are applied. In both benchmarks, the volatility parameter is updated daily according to the prevailing IV surface, which aligns the hedging strategies with dynamic market conditions. The explicit formulas for these two benchmarks are provided in Appendix C.
For all three benchmarks, we further enhance the performance by incorporating the no-trade region, as defined in Equation (4), to ensure a fair and consistent comparison. [^8] The optimization process is carried out as detailed in Section 2.3, following Equation (9), using Mini-batch Stochastic Gradient Descent. Additionally, $l$ is optimized based on the risk measure used to benchmark our framework, with each benchmark having its distinct optimal value for $l$ [^9] For instance the DG strategy has a different threshold $l$ depending on which risk measure is used for evaluation.
@@ 4.3.1   Neural network architecture
We consider a RNN-FNN architecture with two LSTM cells of width 56, two FFNN-hidden layers of width 56 with ReLU activation function (i.e., $g_{\mathcal{L}_{i}}(X)=\text{max}(0,X)$ for $i=1,2$), and one two-dimensional output FFNN layer with a linear activation function. Numerical experiments suggest that the parameter $\lambda$ associated with the soft constraint should be set to one. A detailed description of the experimental procedure which led to such choice can be found in Appendix D.
Agents are trained as described in Section 2.3 on a training set of 400,000 independent simulated paths with mini-batch size of 1000 and an initial learning rate of 0.0005. In addition, we include dropout regularization method with parameter $p=0.5$ as in François et al., (2024). The training procedure is implemented in Python, using Tensorflow and considering the Glorot and Bengio, (2010) random initialization of the initial parameters of the neural network. Numerical results are obtained from a test set of 100,000 independent paths.
@@ 4.3.2   State space
The state space considered in our RL framework includes the state variables generated by the JIVR model, along with a new set of state variables associated with the straddle and hedging portfolio. These variables are detailed in Table 1.
@@
$\tau_{t}$ Time-to-maturity of the straddle
$\{\beta_{t,i}\}^{5}_{i=1}$ IV factors described in Section 3.1
@@
  • For all Black-Scholes Greeks, as well as the portfolio value and hedging option value, we use the implied volatility $\sigma(M_{t},\tau_{t},\beta_{t})$ from the static surface as the volatility input parameter.
In our illustrative example, the RL agent seeks to hedge a straddle contract with the same specifications across different market dynamics. According to the terminology of Peng et al., (2024), this problem is a contract-specific reinforcement learning task, where the optimization problem is solved for a given contract with predefined parameters. Variables related to the target portfolio (such as $\mathcal{P}_{t}$, $\Delta_{t}^{P}$, and $\Gamma_{t}^{P}$) are not strictly necessary, as they can theoretically be recovered by the ANN if needed. However, our numerical experiments demonstrate that in practice their inclusion enhances training performance across all risk measures (details in Appendix E). Furthermore, incorporating these state variables extends our framework to enable its application in a contract-unified setting, allowing for the optimization of portfolios with any combination of options and contract parameters.
@@ 4.4.1   Benchmarking in the absence of transaction costs
$\text{mean}\left(\xi_{T}^{(\tilde{\phi}_{\theta},\,l)}\right)=\frac{1}{N}\sum_{i=1}^{N}\xi_{T,i}^{(\tilde{\phi}_{\theta},\,l)},$
@@
where $\xi_{T,i}^{(\tilde{\phi}_{\theta},\,l)}$ represents the $i$-th terminal hedging error in the test set of size $N$. Additionally, we incorporate the sample standard deviation of the terminal hedging error, $\text{std}\left(\xi_{T}^{(\tilde{\phi}_{\theta},\,l)}\right)$, as a metric to quantify the variability of hedging errors within the test set. Our analysis is conducted under the assumption of zero transaction costs, i.e., $\kappa_{1}=\kappa_{2}=0$.
@@
Table 2: Hedging performance metrics under the assumption of zero transaction costs.
Instruments $S_{t}$ $S_{t}$+$\text{O}_{t}$
Strategy D RL DG RL
MSE SMSE CVaR95% MSE SMSE CVaR95%
$\text{mean}\left(\xi_{T}^{(\tilde{\phi}_{\theta},\,l)}\right)$0.713 -0.543 -0.656 -0.681 -0.069 -0.035 -0.089 -0.087
$\text{std}\left(\xi_{T}^{(\tilde{\phi}_{\theta},\,l)}\right)$1.756 -1.392 -1.526 -1.702 -0.811 -0.324 -0.325 -0.326
MSE -3.593 -2.232 -2.760 -3.362 -0.663 -0.106 -0.114 -0.114
SMSE -1.193 -0.546 -0.424 -0.596 -0.338 -0.038 -0.025 -0.027
CVaR95% -3.606 -2.549 -2.208 -2.031 -1.927 -0.648 -0.516 -0.514
@@
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa_{1}=\kappa_{2}=0$). Agents are trained according to the conditions outlined in Section 4.3. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. Columns under $S_{t}$ represent hedging with the risk-free and underlying assets, while those under $S_{t}+O_{t}$ incorporate an ATM call option with an initial maturity of $T^{*}=84$ days. D stands for delta hedging, whereas DG refers to delta gamma hedging.
@@
Table 2 presents the optimal values of the risk measures for the various hedging strategies in two cases. In the first case, the hedging instruments are limited to the risk-free asset and the underlying asset (columns labeled as $S_{t}$). In the second scenario, the ATM call option is introduced as an additional hedging instrument (columns labeled as $S_{t}+\text{O}_{t}$). The columns under RL correspond to different risk measures used as objective functions during training, while each row represents the performance metric computed from test set hedging errors. In both cases, RL strategies consistently outperform the benchmarks and achieve the optimal values when the performance assessment metric matches the risk measure used during training.
Moreover, our numerical results highlight the benefits of incorporating a second hedging instrument. Specifically, all strategies that include an option as an additional hedging instrument exhibit lower risk in terms of standard deviation, MSE, SMSE, and CVaR95%, compared to those relying solely on a single hedging instrument, including RL-based strategies. Notably, the delta-gamma hedging strategy reduces standard deviation by at least 42%, as seen by comparing the lowest standard deviation among single-instrument strategies (1.392) to that of the DG strategy (0.811). It also achieves a 70% reduction in MSE, a 20% decrease in SMSE, and a 5% reduction in CVaR relative to the lowest values of these performance metrics across all strategies in the $S_{t}$ column.
RL agents utilizing multiple hedging instruments achieve significantly lower risk than the DG strategy, with a standard deviation reduction of at least 60%—comparing the highest standard deviation among RL strategies with two hedging instruments (0.326) to that of the DG strategy (0.811). This advantage extends to other performance metrics, with reductions of at least 92% in CVaR95% and 73% in SMSE.
Figure 1 depicts the distribution of hedging errors across various strategies. Panel A contrasts the hedging error distributions of the benchmark and RL agents—both using only the underlying asset—with the traditional DG strategy, showing that incorporating an option significantly reduces risk. Panel B compares the DG strategy to the RL-MSE strategy, both utilizing three hedging instruments, highlighting the RL approach’s superior performance in variance reduction, consistently with Table 2. Finally, Panel C compares the three RL agents, revealing that strategies based on asymmetric risk measures produce distributions with greater skewness.
@@
  • Results are computed using 100,000 out-of-sample paths according to the conditions outlined. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. RL strategies labeled as "one" represent hedging with the risk-free and underlying assets; those labeled as "two" incorporate an ATM call option with a maturity of $T^{*}=84$ days.
@@ 4.4.2   Benchmarking in the presence of transaction costs
This analysis incorporates the no-trade region, defined by Equation (4), to optimize rebalancing frequency while accounting for transaction costs. For benchmarks, the rebalancing threshold $l$ is estimated using Equation (9) on the training set, treating each combination of risk measures and transaction cost levels as independent optimization problems. In contrast, RL strategies estimate this parameter jointly with other ANN parameters during training. Table 3 reports the optimal rebalancing thresholds $l$ across different transaction cost levels for both DG and RL strategies, considering all risk measures.
Table 3: Optimal rebalancing threshold $l$ values for DG and RL strategies.
Risk measure MSE SMSE CVaR95%
$\kappa_{1}$ $\kappa_{2}$ DGl RLl DGl RLl DGl RLl
$0\%$ $0\%$ 0.0--- 0.0--- 0.0--- 0.0--- 0.0--- 0.0---
$0.05\%$ $0.5\%$ 0.904 0.013 1.777 0.019 2.011 0.018
$0.05\%$ $1\%$ 1.107 0.017 2.425 0.023 2.520 0.026
$0.05\%$ $1.5\%$ 1.205 0.032 2.502 0.032 2.671 0.033
$0.05\%$ $2\%$ 1.498 0.033 2.517 0.034 2.736 0.034
The numerical results presented in Table 3 show a monotonic increase in rebalancing threshold values as transaction costs rise, a pattern consistently observed across all risk measures for both benchmarks and RL agents. This trend reflects broader no-trade regions at higher transaction cost levels, suggesting that small adjustments increasingly degrade hedging performance as transaction costs increase, regardless of the risk measure or approach. The incorporation of no-trade regions proves beneficial, as evidenced by the non-zero rebalancing threshold values when transaction costs are introduced into the hedging problem. In contrast, when transaction costs are absent, the zero threshold values obtained from the optimization process are expected. This is because, under these conditions, rebalancing does not negatively affect hedging performance.
The optimal rebalancing thresholds for RL agents are consistently and substantially lower than those of non-RL strategies (columns 2, 4, and 6), often nearing zero across all transaction cost levels. This suggests that the no-trade region serves as a noise-reduction mechanism within the RL framework. Indeed, the continuity of the ANN function makes identical hedging positions at consecutive time steps unlikely, leading to frequent small adjustments that accumulate transaction costs and reduce performance. The no-trade region mitigates this issue by preventing unnecessary trades.
Table 4: Optimal risk measure values of deep hedging, delta hedging, and delta-gamma hedging.
Risk measure MSE SMSE CVaR95%
Instruments $S_{t}$ $S_{t}$+$\text{O}_{t}$ $S_{t}$ $S_{t}$+$\text{O}_{t}$ $S_{t}$ $S_{t}$+$\text{O}_{t}$
$\kappa_{1}$ $\kappa_{2}$ Dl RLl DGl RLl Dl RLl DGl RLl Dl RLl DGl RLl
Panel A ($l=0$)
$0\%$ $0\%$ 3.593 2.232 0.663 0.106 1.193 0.424 0.338 0.025 3.606 2.031 1.927 0.514
$0.05\%$ $0.5\%$ 3.384 2.145 0.837 0.124 1.395 0.693 0.723 0.058 3.880 2.281 2.581 0.704
$0.05\%$ $1\%$ 1.092 0.144 1.018 0.099 2.857 0.784
$0.05\%$ $1.5\%$ 1.465 0.165 1.414 0.132 3.159 0.952
$0.05\%$ $2\%$ 1.957 0.193 1.919 0.151 3.487 1.010
Panel B ($l\neq 0$)
$0\%$ $0\%$ 3.593 2.232 0.663 0.106 1.193 0.424 0.338 0.025 3.606 2.031 1.927 0.514
$0.05\%$ $0.5\%$ 3.383 2.124 0.711 0.122 1.361 0.689 0.490 0.052 3.842 2.278 1.935 0.647
$0.05\%$ $1\%$ 0.821 0.136 0.616 0.069 2.015 0.733
$0.05\%$ $1.5\%$ 0.986 0.156 0.803 0.098 2.213 0.863
$0.05\%$ $2\%$ 1.197 0.179 1.025 0.141 2.428 0.972
@@
  • Optimal values of risk measures are computed using 100,000 out-of-sample paths. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. Columns under $S_{t}$ represent hedging with the risk-free and underlying assets, while those under $S_{t}+O_{t}$ incorporate an ATM call option with an initial maturity of $T^{*}=84$ days.
Table 4 presents the optimal risk measures values for two hedging setups: one using only the risk-free asset and the underlying asset (columns $S_{t}$), and another incorporating an ATM call option as an additional hedging instrument (columns $S_{t}+\text{O}_{t}$). The comparison is structured across two panels: Panel A examines strategies without a no-trade region ($l=0$), while Panel B includes it, highlighting its impact on hedging performance.
The impact of the no-trade region is evaluated by comparing each strategy’s performance between Panel A and Panel B. For benchmark strategies, our results show that incorporating a no-trade region significantly improves DG hedging performance across all risk measures, especially as transaction costs increase. For instance, when optimizing the rebalancing threshold using MSE, the optimal MSE for DG strategies decreases by 15%, from 0.837 (Panel A) to 0.711 (Panel B), when the hedging option’s transaction cost parameter is $\kappa_{2}=0.5\%$. This improvement becomes even more pronounced at $\kappa_{2}=2\%$, where MSE drops by 38%, from 1.957 to 1.197, after introducing the no-trade region.
For delta hedging, the benefits of introducing a rebalancing threshold are minimal, given the negligible transaction costs on the underlying asset. Likewise, for RL agents, performance gains from the no-trade region are less pronounced, reinforcing its role as a regularization mechanism that removes very small rebalancing actions. This aligns with the consistently low rebalancing threshold values reported in Table 3.
Given the favorable impact of the no-trade region on hedging performance, we focus on analyzing Panel B, which incorporates this feature. The results show that RL agents consistently outperform benchmarks across all risk measures and choices of hedging instruments. Using the MSE as both the training objective and the performance metric, adding an ATM call option significantly improves hedging performance as observed earlier. In particular, this holds for the DG strategy, which outperforms RL agents not using options. However, this advantage disappears for asymmetric risk measures when hedging option transaction costs exceed 1.5% for SMSE and 2% for CVaR. In such cases, DG strategies fail to outperform RL agents using only a single risky instrument.
The benefits of incorporating a hedging option are particularly evident for RL agents. As shown in the $S_{t}+\text{O}_{t}$ columns, RL agents consistently outperform all benchmarks across various transaction cost levels and risk measures, as indicated by the bold values in each risk measure column. For example, when transaction costs are set to zero, the RL agent trained with MSE using three hedging instruments (column $S_{t}+\text{O}_{t}$ under MSE) and optimized achieves an MSE of 0.106. In contrast, benchmark strategies yield significantly higher MSE values: 3.593 for delta hedging, 2.232 for RL strategies with two hedging instruments, and 0.663 for delta-gamma hedging. This advantage becomes even more pronounced as transaction costs rise, with a similar trend observed across all risk measures.
@@
  • Results are computed using 100,000 out-of-sample paths according to the conditions outlined in Section 4.3. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. The transaction cost parameter for the underlying asset is set to $\kappa_{1}=0.05\%$.
To further highlight the advantages of RL over DG, Figure 2 presents histograms of hedging error distributions at maturity for both strategies under two different transaction cost scenarios. As shown, RL agents consistently produce narrower distributions across all risk measures, indicating greater resilience to rising transaction costs. This stability is particularly beneficial from a risk management perspective, as it ensures more reliable performance despite increasing costs.
4.4.3   Impact of no-trade regions
The hedging cost
$\mbox{HC}_{l}=\sum_{t=0}^{T-1}e^{-r\Delta t}\mathcal{HC}_{t},$ (12)
  • Results are computed over 100,000 out-of-sample paths according to the conditions outlined in Section 4.3.
Conversely, as $\kappa_{2}$ increases, RL agents retain high rebalancing frequency, but keep average transaction costs to a level similar to DG. Thus, more gradual and frequent adjustments from RL mitigate risk more effectively than DG as documented in Section 4.4.2, while leading to similar transaction costs.

### v3

@@ 3.1  Daily implied volatility surfaces
The time-$t$ IV associated to an option with time-to-maturity $\tau_{t}=\frac{T-t}{252}$ years and (scaled) moneyness $M_{t}=\frac{1}{\sqrt{\tau_{t}}}\log\frac{S_{t}e^{(r_{t}-q_{t})\tau_{t}}}{K}$ is modeled as
@@
The vector $\beta_{t}=(\beta_{t,1},\beta_{t,2},\beta_{t,3},\beta_{t,4},\beta_{t,5})$ represents the IV factor coefficients at time $t$, while the functions $\{f_{i}\}_{i=1}^{5}$ allow representing the long-term at-the-money (ATM) level, the time-to-maturity slope, the moneyness slope, the smile attenuation, and the smirk, respectively. A detailed description of the functional components $\{f_{i}\}_{i=1}^{5}$ of the IV surface can be found in Section˜C.1.
@@ 3.2  Joint implied volatility and return
The JIVR model introduced by François et al., (2023) builds upon the IV representation (8), offering an explicit formulation for the joint dynamics of the IV surface and the S&P 500 price. This joint representation is based on an econometric model for (i) the underlying asset returns, and (ii) fluctuations of the IV surface coefficients $\beta_{t}$ along with a mean-reversion component for their volatilities $h_{t}$. The multivariate time series of the JIVR model is provided in Section˜C.2.
The JIVR model is used to generate paths of the state variables $(S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$, which drive the market dynamics, where $h_{t,R}$ and $\{h_{t,i}\}^{5}_{i=1}$ are volatilities for the S&P 500 and each of the IV factors. Estimates of the model parameters and volatility series $\{\hat{h}_{t,i}\}_{t=1}^{N}$ with $i\in\{1,\ldots,5,R\}$ are taken from François et al., (2023).444 François et al., (2023) use a maximum likelihood approach on a multivariate time series made of S&P 500 returns and surface coefficients estimates $\{\hat{\beta}_{t}\}_{t=1}^{N}$, with sample dates extending between January 4, 1996 and December 31, 2020.
@@ 4.1  Market settings for numerical experiments
We consider daily trading periods. For each simulated path, initial conditions of the JIVR model, $\left(\{\beta_{0,i}\}^{5}_{i=1},h_{0,R},\{h_{0,i}\}^{5}_{i=1}\right)$, are randomly sampled from the daily estimated values in our data set, covering the period from January 4, 1996, to December 31, 2020. Across all experiments, the annualized continuously compounded risk-free rate and dividend yield are assumed to remain constant, with values fixed at $r=2.66\%$ and $q=1.77\%$, respectively. [^5] The annualized rates of the S&P 500 dividend yield (1.77%) and the zero-coupon yield (2.66%) are calculated as the average over the sample period from January 4, 1996, to December 31, 2020, using OptionMetrics data. Without loss of generality, the initial value of the underlying asset is set to $S_{0}=100$.666 In our setting, the value of the portfolio to be hedged is proportional to the underlying asset initial value. The hedged portfolio is an ATM straddle with a maturity of $T=63$ days. At any time $t<T$, the portfolio value $\mathcal{P}_{t}$ is determined using the IV surface prevailing at that moment.
The hedging instruments are the risk-free asset, the underlying asset, and an option with a maturity longer than that of the straddle—specifically, an ATM European call option with an initial maturity of $T^{*}=84$ days. Positions in all hedging instruments are rebalanced daily.
The hedge follows the self-financing dynamics from Equation (1), incorporating proportional transaction costs on both the underlying asset and the hedging option. As reported in Chaudhury, (2019), the average cost for S&P 500 index call options is 0.95%. To evaluate its impact, we consider $\kappa_{2}\in\{0.5\%,1\%,1.5\%,2\%\}$. In contrast, transaction costs for the underlying asset are negligible, around 0.047% according to Bazzana and Collini, (2020). We set $\kappa_{1}=0.05\%$. The initial hedging portfolio value matches the straddle price, i.e., $V_{0}=P_{0}$.
@@ 4.2  Benchmarks
For the second and third benchmarks, the delta and gamma of financial instruments are computed using the practitioner’s approach, i.e., using the current IV value. In the case of delta hedging, the delta is adjusted based on the correction introduced by Leland, (1985), which accounts for the impact of proportional transaction costs on the underlying asset position. In both benchmarks, the volatility parameter is updated daily according to the prevailing IV surface, which aligns the hedging strategies with dynamic market conditions. The explicit formulas for these two benchmarks are provided in Appendix˜D.
For all three benchmarks, we further enhance the performance by incorporating the no-trade region, as defined in Equation (10). [^7] The optimization process is carried out as detailed in Section 2.3, following Equation (7), using Mini-batch Stochastic Gradient Descent. Additionally, the no-trade boundary $\ell$ is optimized separately for each risk measure used for benchmarking, with each benchmark exhibiting its own distinct optimal value of $\ell$. Further details are provided in Section˜D.3.
@@ 4.3.1 Neural network architecture
We consider a RNN-FNN architecture with two LSTM cells of width 56, two FFNN-hidden layers of width 56 with ReLU activation function (i.e., $g_{\mathcal{L}_{i}}(X)=\text{max}(0,X)$ for $i=1,2$), and one two-dimensional output FFNN layer with a linear activation function. Numerical experiments detailed in Appendix˜J from the Supplementary Material suggest the value $\lambda=1$ for the soft constraint hyperparameter, which is learned from the validation set.
Agents are trained as described in Section˜2.3 on a training set of 400,000 independent simulated paths with mini-batch size of 1000 and an initial learning rate of 0.0005. In addition, we include dropout regularization method with parameter $p=0.5$ as in François et al., (2024). The training procedure is implemented in Python, using Tensorflow and considering the Glorot and Bengio, (2010) random initialization of the initial parameters of the neural network. The performance assessment is obtained from a test set of 100,000 independent paths.
@@ 4.3.2 State space
The state space presented in Table 1 includes the state variables generated by the JIVR model, along with a new set of state variables associated with the straddle and hedging portfolio.
@@
$\{\beta_{t,i}\}^{5}_{i=1}$ IV factors described in Section˜3.1
@@
  • For all Greeks, as well as the portfolio value and hedging option value, we use the implied volatility $\sigma(M_{t},\tau_{t},\beta_{t})$ from the static surface as the volatility input parameter.
In our illustrative example, the RL agent seeks to hedge a straddle contract with the same specifications across different market dynamics. According to the terminology of Peng et al., (2024), this problem is a contract-specific reinforcement learning task, where the optimization problem is solved for a given contract with predefined parameters. Variables related to the target portfolio (such as $\mathcal{P}_{t}$, $\Delta_{t}^{P}$, and $\Gamma_{t}^{P}$) are not strictly necessary, as they can theoretically be recovered by the ANN if needed. However, our numerical experiments demonstrate that in practice their inclusion enhances training performance across all risk measures (details in Appendix˜E). Furthermore, incorporating these state variables extends our framework to enable its application in a contract-unified setting, allowing for the optimization of portfolios with any combination of options and contract parameters.
@@ 4.4.1 Benchmarking in the absence of transaction costs
$\text{mean}\left(\xi_{T}^{\tilde{\phi}_{\theta}}\right)=\frac{1}{N}\sum_{i=1}^{N}\xi_{T,i}^{\tilde{\phi}_{\theta}},$
@@
where $\xi_{T,i}^{\tilde{\phi}_{\theta}}$ represents the $i$-th terminal hedging error in the test set of size $N$. Additionally, we incorporate the sample standard deviation of the terminal hedging error, $\text{std}\left(\xi_{T}^{\tilde{\phi}_{\theta}}\right)$, as a metric to quantify the variability of hedging errors within the test set. Our analysis is conducted under the assumption of zero transaction costs, i.e., $\kappa_{1}=\kappa_{2}=0$.
@@
Figure 1: Hedging performance metrics under the assumption of zero transaction costs.
@@
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa_{1}=\kappa_{2}=0$). Agents are trained according to the conditions outlined in Section˜4.3. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55 across all initial conditions. Methods denoted by (S) represent hedging with the risk-free and underlying assets, while those denoted by (S + O) incorporate an ATM call option with an initial maturity of $T^{*}=84$ days. D stands for delta hedging, DG denotes delta-gamma hedging, and RL refers to reinforcement learning strategies.
@@
Figure 1 presents the risk measures for the various hedging strategies in two cases. In the first case (the first four columns for each metric), the hedging instruments are limited to the risk-free asset and the underlying asset. In the second scenario (the last four columns), the ATM call option is introduced as an additional hedging instrument. In both cases, RL strategies consistently outperform the benchmarks and achieve the optimal values when the performance assessment metric matches the risk measure used during training. Our numerical results highlight the benefits of incorporating a second hedging instrument. Specifically, all strategies that include an option as an additional hedging instrument exhibit lower risk in terms of standard deviation, MSE, SMSE, and $\text{CVaR}_{95\%}$, compared to those relying solely on a single hedging instrument. Notably, for tail risk captured by CVaR95%, the RL agent—trained on the money account and the underlying asset only—achieves a performance comparable to that of delta-gamma hedging.
@@
  • Results are computed using 100,000 out-of-sample paths. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. RL strategies labeled as "One" represent hedging with the risk-free and underlying assets; those labeled as "Two" incorporate an ATM call option with an initial maturity of $T^{*}=84$ days.
Figure 2 depicts the distribution of hedging errors across various strategies. Panel A contrasts the hedging error distributions of the benchmark and RL agents—both using only the underlying asset—with the traditional DG strategy, showing that incorporating an option significantly reduces risk. Panel B compares the DG strategy to the RL-MSE strategy, both utilizing three hedging instruments, highlighting the RL approach’s superior performance in variance reduction. Finally, Panel C compares the three RL agents, revealing that strategies based on asymmetric risk measures produce distributions with greater skewness.
@@ 4.4.2 Benchmarking in the presence of transaction costs
We now measure the impact of transaction costs on the hedging performance.
Figure 3 displays the optimal values of risk measures for two distinct hedging configurations: one relying solely on the risk-free asset and the underlying asset (first four columns in each group), and another that includes an ATM call option as an additional hedging instrument (last four columns). The comparison contrasts strategies without a no-trade region (solid bars) against those incorporating a no-trade region (striped bars). The no-trade region primarily benefits delta-gamma hedging. In the case of delta hedging, transaction costs associated with trading the underlying asset are minimal and have negligible impact on performance. For reinforcement learning (RL) approaches, trading costs are already internalized within the optimization of the neural network policy, rendering the additional constraint of a no-trade region unnecessary. This is consistent with the persistently low threshold values reported in Figure 13 of Appendix˜A.
Figure 3: Hedging performance in the presence of transaction costs.
MSE Refer to caption
SMSE Refer to caption
CVaR Refer to caption
@@
  • Performance metrics are computed using 100,000 out-of-sample paths. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. Methods denoted by (S) represent hedging with the risk-free and underlying assets, while those denoted by (S + O) incorporate an ATM call option with an initial maturity of $T^{*}=84$ days. D stands for delta hedging, DG denotes delta-gamma hedging, and RL refers to reinforcement learning strategies. Striped bars represent strategies that include the no-trade region, while plain bars correspond to those without it.
RL agents consistently outperform benchmarks across all risk measures and choices of hedging instruments. Using the MSE as a performance metric, adding an ATM call option as a hedging instrument significantly improves hedging performance. In particular, this holds for the DG strategy, which outperforms RL agents not using options. In terms of downside risk management—assessed via SMSE and CVaR metrics—it is noteworthy that, in the presence of transaction costs, the RL algorithm that relies solely on the risk-free asset and the underlying asset either outperforms or provides a performance similar to that of delta-gamma hedging. RL algorithms that incorporate an option as part of the hedging instruments achieve even stronger performance.
@@
  • Results are computed using 100,000 out-of-sample paths according to the conditions outlined in Section˜4.3. The hedged position is an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. The transaction cost parameter for the underlying asset is set to $\kappa_{1}=0.05\%$.
To further highlight the advantage of RL over DG, Figure 4 presents histograms of hedging error distributions at maturity for both strategies under two different transaction cost scenarios. RL agents constantly produce narrower distributions across all risk measures, indicating greater resilience to rising transaction costs. This stability is particularly beneficial from a risk management perspective, as it ensures more reliable performance despite increasing costs.